# Self-supervised 3D Point Cloud Completion via Multi-view Adversarial Learning

### v2

@@ Abstract
In real-world scenarios, scanned point clouds are often incomplete due to occlusion issues. The task of self-supervised point cloud completion involves reconstructing missing regions of these incomplete objects without the supervision of complete ground truth. Current self-supervised methods either rely on multiple views of partial observations for supervision or overlook the intrinsic geometric similarity that can be identified and utilized from the given partial point clouds. In this paper, we propose MAL-SPC, a framework that effectively leverages both region-level and category-specific geometric similarities to complete missing structures. Our MAL-SPC does not require any 3D complete supervision and only necessitates a single partial point cloud for each object. Specifically, we first introduce a Pattern Retrieval Network to retrieve similar position and curvature patterns between the partial input and the predicted shape, then leverage these similarities to densify and refine the reconstructed results. Additionally, we render the reconstructed complete shape into multi-view depth maps and design an adversarial learning module to learn the geometry of the target shape from category-specific single-view depth images. To achieve anisotropic rendering, we design a density-aware radius estimation algorithm to improve the quality of the rendered images. Our MAL-SPC yields the best results compared to current state-of-the-art methods. We will make the source code publicly available at https://github.com/ltwu6/malspc.
@@
Point cloud completion, shape completion, self-supervised learning, generative adversarial learning
@@ I Introduction
To solve this problem, point cloud completion methods have emerged. Point cloud completion aims to infer the missing regions of an object based on the partial point cloud. Most of the existing point cloud completion methods rely on a large number of complete point clouds as supervision [15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. However, in real-world scenes, clean and complete point clouds are really few since current devices struggle to directly capture such data.
@@
To overcome these limitations, some research has focused on unsupervised learning approaches for point cloud completion. Among these methods, some employ complete shapes unpaired with the partial input as priors to drive the completion process [25, 26, 27, 28, 29, 30, 31], while others utilize 2D signals such as RGB images or silhouette maps from multiple viewpoints as supervision [32, 33]. However, the former still requires complete point clouds to acquire shape priors, and the latter relies on multi-view images along with the projection parameters. Recently, self-supervised methods, which only require partial inputs for supervision, have been proposed. These methods estimate the missing parts of the point cloud by utilizing the multi-view consistency of partial scans and the robust learning capabilities of the masked models. However, they face some limitations. First, these methods rely on multi-view partial point clouds of each object as shape constraints or as training data to drive the shape learning process. Nevertheless, obtaining point clouds from multiple viewpoints is also impractical in real-world applications. Secondly, they fail to utilize the inherent local similarities and category-specific geometric properties of objects.
@@
From our perspective, it is commonly observed that an object frequently exhibits similar or repetitive local structures. For instance, a table typically features four identical cylindrical legs. This observation implies that the existing and missing sections of a point cloud are likely to share similar local geometric patterns. Moreover, objects within the same category generally display consistent structures. For example, most airplanes feature two flat wings. Consequently, these inherent geometric similarities, specific to each category, can be utilized to guide the completion of missing regions.
@@
In this paper, we propose MAL-SPC, a Multi-view Adversarial Learning framework for Self-supervised Point cloud Completion. One of the core components, the Pattern Retrieval Network (PRN), reconstructs complete shapes by initially generating a coarse shape and subsequently encoding local geometric patterns from both the coarse and partial input shapes. These patterns manifest in two forms: position encodings, quantified as the average distance between each point and its nearest neighbors, and curvature encodings, defined by the variations in normals between each point and its closest neighbors. Subsequently, a cross-attention-based module is introduced to retrieve correlated local patterns for the coarse shape from the partial shape. These retrieved patterns are further utilized to guide the densification and refinement of the coarse shape, capitalizing on the region-level geometric similarities and repetitiveness inherent in the data.
@@
In addition, we introduce a Multi-view Adversarial Network (MAN) to learn the shared geometry within each category. Objects within the same category commonly exhibit remarkably similar shape distributions because they share many typical geometric structures. This similarity suggests that missing regions of a point cloud are likely to resemble the geometric structure of some objects in the same category. To exploit this similarity, we employ an adversarial learning strategy to learn the distinctive shape characteristics from category-specific data. Specifically, since depth maps provide a more compact and regular shape representation and are easier to acquire than partial point clouds, we render the reconstructed complete point cloud into depth maps from multiple viewpoints to capture the view-specific structures of the shape. Then, we utilize a 2D CNN-based network to discriminate the rendered depth maps and those from the same category in the dataset, thus refining our predictions based on category-specific geometric knowledge. Additionally, to address the anisotropic rendering issue, we propose a density-aware radius estimation algorithm to adaptively adjust the point radius from different viewpoints, thereby improving the quality of rendered images.
@@
Our MAL-SPC is unique in that it does not require multiple views of a partial point cloud for supervision or as additional data to drive the training process. Instead, it can achieve remarkable performance by employing only a single partial point cloud per object for training and supervision. Additionally, MAL-SPC effectively leverages the intrinsic geometry similarity from partial data to reconstruct missing structures. Furthermore, unlike other self-supervised methods, our approach does not require any data preprocessing processes, such as random region removal, which is time-consuming and computationally intensive.
We conduct extensive experiments on both synthetic datasets and real-world scans. The experimental results demonstrate that our model outperforms the state-of-the-art approaches.
@@
  • We propose a Multi-view Adversarial Network that facilitates the completion of missing shapes by leveraging the category-specific geometric similarities. To achieve anisotropic rendering and improve the quality of the rendered depth maps, we design a density-aware radius estimation algorithm.
@@
Figure 1: Flowchart of the proposed self-supervised point cloud completion method. Given a depth map $\mathbf{D}_{0}$, we can obtain an incomplete point cloud $\mathbf{P}_{in}$ by back-projecting it into 3D space. We first employ an auto-encoder to generate a coarse shape, and then introduce a pattern retrieval module to guide the refinement and densification of the shape based on the position and curvature patterns of $\mathbf{P}_{in}$. Then, we render the generated point clouds into silhouette maps from the input viewpoint and depth maps from various random viewpoints. The silhouette maps are supervised by the mask image binarized from $\mathbf{D}_{0}$, while the depth maps are used for adversarial learning with the counterparts from the same category.
@@ II-A Supervised Point Cloud Completion
The voxel-based methods [34, 35, 36, 37, 38, 39, 40] first convert partial point clouds into 3D grids through voxelization and subsequently utilize the 3D Convolutional Neural Networks for learning the complete shapes. Dai et al. [34] employed a 3D-Encoder Predictor Network to generate an initial approximation of a shape, followed by a 3D shape synthesis method that refines the coarse output using a shape prior derived from a shape database. Xie et al. [35] introduced a Gridding module along with a reverse module for seamless conversion between point clouds and voxels, thereby minimizing structural information loss during the voxelization process. Additionally, they devised a Cubic Feature Sampling module to extract contextual features from neighboring points. Sun et al. [36] represented point clouds as occupancy grids and leveraged local patches of partial inputs to infer missing parts by capitalizing on the potential repetition of local patterns within an object. Initially, they employed a network to capture similar local patches of the inputs. Then, these patches undergo deformation and blending processes, ultimately reconstructing into the final outputs. Voxel-based methods can utilize the powerful feature extraction capabilities of CNNs. However, the computation speed of 3D CNN is relatively slow, and the conversion process between point clouds and voxels inevitably introduces accuracy loss.
The point-based methods [15, 41, 42, 43, 44, 45, 46] directly encode and decode features within the 3D point space by using Multi-Layer Perceptron (MLP). To obtain smooth and detailed surfaces, some efforts employed the Folding operation, originally proposed in FoldingNet [41], to generate complete point clouds. Yuan et al. [15] designed an MLP-based encoder to extract the global feature based on the partial input. Then, they leveraged fully connected (FC) layers to reconstruct a coarse point cloud. Subsequently, the Folding operation was applied to each coarse point. This approach effectively capitalizes on the inherent flexibility of FC layers and the desirable surface-smoothing characteristics offered by the Folding operation. Groueix et al. [47] employed multiple MLPs to deform 2D grids into distinct local patches, enabling the generation of more intricate shapes compared to FoldingNet. However, this approach resulted in overlapping regions between patches. To address this issue, Liu et al. [48] proposed an expansion penalty method to minimize the overlap areas. Besides, it is difficult for a fixed 2D grid to capture the intricate topologies of point clouds. To improve the shape representation ability of the seeds, Yang et al. [49] design a model for learning high-dimensional seeds for generating complete shapes. Likewise, Pang et al. [50] utilized a network to learn the point cloud topology and tore the grid into patches to better represent the object’s shape.
@@ II-B Unsupervised Point Cloud Completion
Due to the challenges in acquiring complete point clouds in real-world scenarios, some studies have proposed unsupervised methods for point cloud completion.
@@
Unsupervised point cloud methods do not require paired complete 3D ground truth to drive the completion process. These methods can be divided into three categories according to the supervision signal, i.e., unpaired methods, 2D-supervised methods, and self-supervised methods. The unpaired point cloud completion methods generate complete point clouds without paired ground truth as supervision during the training process [25, 26, 27, 28, 29, 30, 31, 29, 62, 63]. In fact, these methods leverage complete and clean point clouds from other domains to train a model to capture 3D shape priors. Then, they used the learned shape priors to guide the completion of the partial point cloud from the source domain. The first unpaired method is PCL2PCL, proposed by Chen et al. [25]. PCL2PCL first used partial point clouds from the source domain to pre-train a partial-shape reconstruction Auto-encoder and then trained a complete-shape reconstruction Auto-encoder by utilizing complete point clouds from another domain. Subsequently, they employed a GAN-based model [64] to transform the partial global features to the same distribution with the complete shapes’ global features. To enhance the learning ability of the generation network, Wen et al. [26] proposed a cycle-based model to learn the transformation between partial and complete point clouds. Cui et al. [63] proposed a latent-space energy-based model to learn a conditional distribution of the complete shape encoding given the partial encoding. Then, they designed a residual sampling strategy to select a valid encoding and decode it into a complete shape. Cao et al. [29] proposed a knowledge transfer model to learn the shape information of partial inputs. They designed a teacher-assistant-student network to transfer the knowledge learned from complete shapes to partial shapes. In addition to the shape priors provided by complete point clouds, Ma et al. [30] proposed to leverage the symmetry of objects to reconstruct the missing regions.
@@
The 2D-supervised methods employ 2D images along with projection parameters to guide the completion of 3D shapes. Wu et al. [32] projected reconstructed point clouds into the 2D plane and then used 2D points within the silhouette maps to supervise the view-specific shape. They also proposed a view-based calibrator to reduce noises in the predicted complete shapes. Aiello et al. [33] enhanced the 3D point clouds’ features with 2D RGB images by attention modules. They introduced a framework that rendered predicted point clouds from certain view angles and used 2D silhouette maps to supervise the renderer images.
@@
The self-supervised methods only leverage the partial point cloud as a supervision signal to guide the completion process. Mittal et al. [65] proposed the first learning-based self-supervised method, PointPnCNet. PointPnCNet first randomly removed a small region from the input partial point cloud and used a network to predict the missing regions, including the manually removed parts and naturally missing parts. It employed the partial matching loss to preserve the existing parts in the predicted shapes. However, PointPnCNet relies on multiple partial observations for supervision, whereas our approach requires only a single-view partial point cloud. Inspired by closed-loop systems in control theory, Hong et al. [66] first predicted a coarse point cloud and used it to synthesize partial point clouds, and then input them to the same model as the initial input. By utilizing consistency shape loss among predictions generated by multiple partial inputs, the model could generate complete shapes. Nevertheless, this method requires the same training on the testing objects. Cui et al. [67] divided partial point clouds into three groups and then proposed a masked autoencoder to reconstruct complete shapes. Rather than directly using partial Chamfer Distance loss, they designed a Region-aware Chamfer Distance loss to reduce the noises of the predictions. Kim et al. [68] proposed a pose-aware approach for self-supervised point cloud completion. They projected the shape features onto various viewpoints and reconstructed the corresponding incomplete point clouds. They leveraged the shape consistency among multiple incomplete point clouds of the same object to constrain the reconstruction process.

### v3

# Unsupervised 3D Point Cloud Completion via Multi-view Adversarial Learning

@@ Abstract
In real-world scenarios, scanned point clouds are often incomplete due to occlusion issues. The tasks of self-supervised and weakly-supervised point cloud completion involve reconstructing missing regions of these incomplete objects without the supervision of complete ground truth. Current methods either rely on multiple views of partial observations for supervision or overlook the intrinsic geometric similarity that can be identified and utilized from the given partial point clouds. In this paper, we propose MAL-UPC, a framework that effectively leverages both region-level and category-specific geometric similarities to complete missing structures. Our MAL-UPC does not require any 3D complete supervision and only necessitates single-view partial observations in the training set. Specifically, we first introduce a Pattern Retrieval Network to retrieve similar position and curvature patterns between the partial input and the predicted shape, then leverage these similarities to densify and refine the reconstructed results. Additionally, we render the reconstructed complete shape into multi-view depth maps and design an adversarial learning module to learn the geometry of the target shape from category-specific single-view depth images of the partial point clouds in the training set. To achieve anisotropic rendering, we design a density-aware radius estimation algorithm to improve the quality of the rendered images. Our MAL-UPC outperforms current state-of-the-art self-supervised methods and even some unpaired approaches. We will make the source code publicly available at https://github.com/ltwu6/malspc.
@@
Point cloud completion, shape completion, unsupervised learning, generative adversarial learning
@@ I Introduction
To solve this problem, point cloud completion methods have emerged. Point cloud completion aims to infer the missing regions of an object based on the partial observation. Most of the existing methods rely on a large number of complete point clouds as supervision [15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. However, in real-world scenes, clean and complete point clouds are really few since current devices struggle to directly capture such data.
@@
To overcome these limitations, some research has focused on unsupervised learning approaches for point cloud completion. Among these methods, some employ complete shapes unpaired with the partial input as priors to drive the completion process [25, 26, 27, 28, 29, 30, 31], and some utilize 2D signals such as RGB images or silhouette maps from multiple viewpoints as supervision [32, 33]. However, the former still requires complete point clouds to acquire shape priors, and the latter relies on multi-view images along with the projection parameters. Recently, self-supervised methods, which only require partial inputs for supervision, have been proposed. These methods estimate the missing parts of the point cloud by utilizing the multi-view consistency of partial scans and the robust learning capabilities of the masked models. However, they face some limitations. First, most methods rely on multi-view partial point clouds of each object as shape constraints or as training data to drive the shape learning process. Nevertheless, obtaining point clouds from multiple viewpoints is also impractical in real-world applications. Second, they fail to utilize the inherent local similarities and category-specific geometric properties of objects.
@@
From our perspective, it is commonly observed that an object frequently exhibits similar or repetitive local structures. For instance, a table typically features four identical cylindrical legs. This observation implies that the existing and missing regions of a point cloud are likely to share similar local geometric patterns. Moreover, objects within the same category generally display consistent structures. For example, most airplanes feature two flat wings. Consequently, these inherent geometric similarities, specific to each category, can be utilized to guide the completion of missing regions.
@@
In this paper, we propose MAL-UPC, a Multi-view Adversarial Learning framework for Unsupervised Point cloud Completion. One of the core components, the Pattern Retrieval Network (PRN), reconstructs complete shapes by initially generating a coarse shape and subsequently encoding local geometric patterns from both the coarse shape and the partial input. These patterns manifest in two forms: position encodings, quantified as the average distance between each point and its nearest neighbors, and curvature encodings, defined by the variations in normals between each point and its closest neighbors. Subsequently, a cross-attention-based module is introduced to retrieve correlated local patterns for the coarse shape from the partial shape. These retrieved patterns are further utilized to guide the densification and refinement of the coarse shape, capitalizing on the region-level geometric similarities and repetitiveness inherent in the data.
@@
In addition, we introduce a Multi-view Adversarial Network (MAN) to learn the shared geometry within each category. Objects within the same category commonly exhibit remarkably similar shape distributions because they share many typical geometric structures. This similarity suggests that missing regions of a point cloud are likely to resemble the geometric structure of other objects in the same category. To exploit this similarity, we employ an adversarial learning strategy to learn the distinctive shape characteristics from category-specific data. Specifically, since depth maps provide a more compact and regular shape representation and are easier to acquire than partial point clouds, we render the reconstructed complete point cloud into depth maps from multiple viewpoints to capture the view-specific structures of the shape. Meanwhile, we construct an image bank for each category using the depth maps from which the partial point clouds in the training set are derived. Finally, we utilize a 2D CNN-based network to discriminate the rendered depth maps and those from the image bank, thus refining our prediction based on category-specific geometric knowledge. Additionally, to achieve anisotropic rendering, we propose a density-aware radius estimation algorithm to adaptively adjust the point radius from different viewpoints, thereby improving the quality of rendered images.
@@
Compared to previous methods, our MAL-UPC offers the following advantages: First, it does not require multi-view partial point clouds for supervision or as additional data to drive the training process. Instead, it can achieve remarkable performance by employing only single-view partial observations, i.e., the 3D (partial point cloud) and 2D (silhouette map) representations of the input for supervision, along with the category-specific single-view depth maps from which the partial point clouds in the training set are derived for adversarial learning. Moreover, MAL-UPC effectively leverages the intrinsic geometry similarity from partial data to reconstruct missing structures. In addition, unlike self-supervised and weakly-supervised methods, our approach does not require any data preprocessing processes, such as random region removal, which is time-consuming and computationally intensive.
We conduct extensive experiments on both synthetic datasets and real-world scans. Since there are no unsupervised 3D point cloud completion methods that only rely on single-view partial observations during training (i.e., our settings), except the self-supervised approaches, the main comparison baselines in this paper are the self-supervised methods. In addition, we also compare our method with some unpaired methods and supervised methods. The experimental results demonstrate that our model outperforms self-supervised methods and even some unpaired approaches. Although our model does not outperform the previous unpaired methods in all instances, it only requires incomplete training data.
@@
  • We propose a Multi-view Adversarial Network that facilitates the completion of missing shapes by leveraging category-specific geometric similarities. To achieve anisotropic rendering and improve the quality of the rendered depth maps, we design a density-aware radius estimation algorithm.
@@
Figure 1: Flowchart of the proposed unsupervised point cloud completion method. Given a depth map $\mathbf{D}_{0}$, we can obtain an incomplete point cloud $\mathbf{P}_{in}$ by back-projecting it into 3D space. We first employ an auto-encoder to generate a coarse shape, and then introduce a pattern retrieval module to guide the refinement and densification of the shape based on the position and curvature patterns of $\mathbf{P}_{in}$. Then, we render the generated point cloud into a silhouette map from the input viewpoint and depth maps from various random viewpoints. The silhouette map is supervised by the mask image binarized from $\mathbf{D}_{0}$, while the depth maps are used for adversarial learning with the category-specific single-view depth maps of the partial point clouds from the training set.
@@ II-A Supervised Point Cloud Completion
The voxel-based methods [34, 35, 36, 37, 38, 39, 40] first convert partial point clouds into 3D grids through voxelization and subsequently utilize the 3D Convolutional Neural Networks for learning the complete shapes. Dai et al. [34] employed a 3D-Encoder Predictor Network to generate an initial approximation of a shape, followed by a 3D shape synthesis method that refines the coarse output using a shape prior derived from a shape database. Xie et al. [35] introduced a Gridding module along with a reverse module for seamless conversion between point clouds and voxels, thereby minimizing structural information loss during the voxelization process. Additionally, they devised a Cubic Feature Sampling module to extract contextual features from neighboring points. Sun et al. [36] represented point clouds as occupancy grids and leveraged local patches of partial inputs to infer missing parts by capitalizing on the potential repetition of local patterns within an object. Initially, they employed a network to capture similar local patches of the inputs. Then, these patches underwent deformation and blending processes, ultimately reconstructing into the final outputs. The Voxel-based methods can utilize the powerful feature extraction capabilities of CNNs. However, the computation speed of 3D CNN is relatively slow, and the conversion process between point clouds and voxels inevitably introduces accuracy loss.
The point-based methods [15, 41, 42, 43, 44, 45, 46] directly encode and decode features within the 3D point space by using the Multi-Layer Perceptron (MLP). To obtain smooth and detailed surfaces, some efforts employed the Folding operation, originally proposed in FoldingNet [41], to generate complete point clouds. Yuan et al. [15] designed an MLP-based encoder to extract the global feature based on the partial input. Then, they leveraged fully connected (FC) layers to reconstruct a coarse point cloud. Subsequently, the Folding operation was applied to each coarse point. This approach effectively capitalizes on the inherent flexibility of FC layers and the desirable surface-smoothing characteristics offered by the Folding operation. Groueix et al. [47] employed multiple MLPs to deform 2D grids into distinct local patches, enabling the generation of more intricate shapes. However, this approach resulted in overlapping regions between patches. To address this issue, Liu et al. [48] proposed an expansion penalty method to minimize the overlap areas. Besides, it is difficult for a fixed 2D grid to capture the intricate topologies of point clouds. To improve the shape representation ability of the seeds, Yang et al. [49] designed a model for learning high-dimensional seeds for generating complete shapes. Likewise, Pang et al. [50] utilized a network to learn the point cloud topology and tore the grid into patches to better represent the object’s shape.
@@ II-B Unsupervised Point Cloud Completion
Due to the challenges in acquiring complete point clouds in real-world scenarios, some studies have proposed unsupervised methods for point cloud completion. Unsupervised point cloud completion methods do not require complete 3D ground truth or partial-complete pairs to drive the completion process. Existing unsupervised methods can be divided into three categories according to the supervision signal, i.e., unpaired methods, 2D-supervised methods, and self-supervised methods.
@@
The unpaired point cloud completion methods generate complete point clouds without paired ground truth as supervision during the training process. Instead, they leverage complete and clean point clouds from other domains to capture 3D shape priors. Then, they used the learned shape priors to guide the completion of the partial point cloud from the source domain [25, 26, 27, 28, 29, 30, 31, 62, 63]. The first unpaired method is PCL2PCL, proposed by Chen et al. [25]. PCL2PCL first used partial point clouds from the source domain to pre-train a partial-shape reconstruction auto-encoder and then trained a complete-shape reconstruction auto-encoder by utilizing complete point clouds from another domain. Subsequently, they employed a GAN-based model [64] to transform the partial global features to the same distribution with the complete shapes’ global features. To enhance the learning ability of the generation network, Wen et al. [26] proposed a cycle-based model to learn the transformation between partial and complete point clouds. Cui et al. [63] proposed a latent-space energy-based model to learn a conditional distribution of the complete shape encoding given the partial encoding. Then, they designed a residual sampling strategy to select a valid encoding and decode it into a complete shape. Cao et al. [29] proposed a knowledge transfer model to learn the shape information of partial inputs. They designed a teacher-assistant-student network to transfer the knowledge learned from complete shapes to partial shapes. In addition to the shape priors provided by complete point clouds, Ma et al. [30] proposed to leverage the symmetry of objects to reconstruct the missing regions. The unpaired methods rely on additional complete point clouds from other domains to learn shape priors. By contrast, our MAL-UPC only requires single-view partial observations in the training set.
@@
The 2D-supervised methods employ 2D images along with projection parameters to guide the completion of 3D shapes. Wu et al. [32] projected reconstructed point clouds into the 2D plane and then used 2D points within the silhouette maps to supervise the view-specific shape. They also proposed a view-based calibrator to reduce noises in the predicted complete shapes. Aiello et al. [33] enhanced the 3D point clouds’ features with 2D RGB images by attention modules. They introduced a framework that rendered predicted point clouds from certain view angles and used 2D silhouette maps to supervise the rendered images. However, the 2D-supervised methods utilize images from multiple viewpoints of an object to provide direct and powerful supervision. Thus, [33] also refers to these methods as weakly-supervised point cloud completion methods. Besides, the existing 2D-supervised methods are also referred to as cross-modal point cloud completion methods because they leverage the features of RGB images to assist in point cloud completion. These configurations differ significantly from the settings of our method. Therefore, the 2D-supervised methods are not included in the comparison baselines of our experiments.
@@
The self-supervised methods only leverage the partial point cloud as a supervision signal to guide the completion process. Mittal et al. [65] proposed the first learning-based self-supervised method, PointPnCNet. PointPnCNet first randomly removed a small region from the input partial point cloud and used a network to predict the missing regions, including the manually removed parts and naturally missing parts. It employed the partial matching loss to preserve the existing parts in the predicted shapes. However, PointPnCNet relies on multiple partial point clouds for supervision, and thus it is also referred to as a weakly-supervised method. By contrast, our approach requires only single-view partial observations. Inspired by closed-loop systems in control theory, Hong et al. [66] first predicted a coarse point cloud and used it to synthesize partial point clouds, and then input them to the same model as the initial input. By utilizing consistency shape loss among predictions generated by multiple partial inputs, the model could generate complete shapes. Nevertheless, this method requires the same training on the testing objects. Cui et al. [67] divided partial point clouds into three groups and then proposed a masked auto-encoder to reconstruct complete shapes. Rather than directly using partial Chamfer Distance loss, they designed a Region-aware Chamfer Distance loss to reduce the noises of the predictions. Kim et al. [68] proposed a pose-aware approach for self-supervised point cloud completion. They projected the shape features onto various viewpoints and reconstructed the corresponding incomplete point clouds. Then, they leveraged the shape consistency among multiple incomplete point clouds of the same object to constrain the reconstruction process.